the second side of a triangle is one centimeter more than twice the first side, and is one centimeters less than the third side. Find the measurements of each if the perimeter of the triangle is 21 centimeters
Answer by 55305(86)

Lets assume that the length of the first side is x. That means the length of the second side is 2x+1, and the length of the third side is 2x+2. So if we add up all the side lengths, we get the perimeter, which is 21.
So x+2x+1+2x+2=21, which simplifies to 5x+3=21. Lets subtract 3 from each side to isolate 5x. 5x+3-3=21-3, 5x=18. Lets divide by 5 to find x, the length of the first side. 5x/5=18/5, x=3.6
So the lengths of the sides are 3.6 cm,8.2 cm, and 9.2 cm.